OpenStudy (anonymous):

The perimeter of a square is 20 units. One side is the square root of a+10. What is the value of a?

myininaya (myininaya):

there are 4 edges that make up a square adding these edges we get 20 each edge has equal length so we have L+L+L+L=20 if L=sqrt(a+10) then what is a you have 4L=20 since (L+L+L+L=4L) just replace L with sqrt{a+10) and solve for a

OpenStudy (anonymous):

I have \[\sqrt{a+10}=5\] \[\sqrt{a+10}^{2}=5^2\] a+10=25 a=15 ?

myininaya (myininaya):

looks good :)